OpenStudy (anonymous):

y=x+5 y=0+5 y=5 this is my y-intercept .. I think ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

How do I find the slope ?

OpenStudy (austinl):

Slope intercept form is, \(y=mx+b\) where, \(m=\text{slope}\) So, what is m in your equation?

OpenStudy (anonymous):

I have no idea .

OpenStudy (anonymous):

keep in mind x is the same as 1*x

OpenStudy (anonymous):

so is it 5=mx+5? im so lost .

OpenStudy (austinl):

\(y=mx+b⇒y=x+5\) \(~~~~~~~~~~~~~~~~~~~~~~y=(1)x+5\)

OpenStudy (anonymous):

not following what I'm suppose to actually DO .

OpenStudy (austinl):

"m" is the slope, now looking at what I posted, what is "m"?

OpenStudy (anonymous):

umm .. *sigh* no .. :I

OpenStudy (austinl):

m=1

OpenStudy (anonymous):

mmhm .. soo y=mx+b my y intercept was 5 so is it 5=1x+b ?

OpenStudy (austinl):

no, b is the y intercept...

OpenStudy (austinl):

\(y=1x+5\)

OpenStudy (anonymous):

oh okay .. so now I'm going to .. make that 6x ?

OpenStudy (austinl):

No.

OpenStudy (anonymous):

:(

OpenStudy (austinl):

\(\huge{\color{red}{\text{Slope Intercept Form}}}\) \(y=mx+b\) \(m=\text{slope}\) \(b=\text{y-intercept}\) \(\Large{\color{blue}{\text{You Have}}}\) \(y=x+5\) \(\rightarrow m=1\) \(\rightarrow b=5\)